OpenStudy (anonymous):

Rewrite y = 3x2 - 24x + 10 in general form.

OpenStudy (mertsj):

General Form You have already seen how to graph a quadratic function when it is in its standard form. Recall that the standard form of a quadratic function looks like: y=a(x-h)^2+k But how do you graph a quadratic function when it is in its general form? The general form of a quadratic function looks like this: y=ax^2+bx+c

OpenStudy (mertsj):

So how does your author define general form?

OpenStudy (anonymous):

h=x, k=y... the problem I'm having with this one is splitting the 3 in 3x^2...?

OpenStudy (anonymous):

Its just asking to rewrite it not graph it... heres what I have so far:

OpenStudy (mertsj):

So you want h-k form?

OpenStudy (anonymous):

-10 on both sides Y-10=3x^2-24x 24/2=12 12^2=144 Y-10=3x^2-24x+144

OpenStudy (anonymous):

Its in standard form, I want general form(:

OpenStudy (mertsj):

Let's write it like this: